### v1

# Violation of local reciprocity in charge–orbital interconversion

Abstract
We demonstrate a violation of local reciprocity in the interconversion between charge and orbital currents. By investigating orbital torque and orbital pumping in W/Ni bilayers, we show that the charge-orbital interconversion in the bulk of the W layer exhibits opposite signs in the direct and inverse processes—the direct and inverse orbital Hall effects being positive and negative, respectively. This finding provides direct evidence of local non-reciprocity in the charge-orbital interconversion, in agreement with a theoretical prediction. These results highlight the unique characteristics of charge-orbital coupled transport and offer fundamental insights into the mechanisms underlying orbital-current-driven phenomena.
Since the discovery of the spin Hall effect (SHE), the interconversion between charge and spin currents through spin-orbit coupling has played a crucial role in the development of spintronics RevModPhys.87.1213 . The direct process of the SHE converts a charge current into a spin current, enabling the electric control of magnetization through spin-orbit torques AndoPRL ; liu2012spin . This process has attracted extensive attention for its topological and quantum mechanical nature, as well as its potential for spintronic applications, such as non-volatile memories and neuromorphic computing RevModPhys.91.035004 . The inverse process of the SHE, the conversion from a spin current into a charge current, offers a way for the electric detection of spin currents saitoh2006conversion ; valenzuela2006direct ; KimuraPRL . This mechanism has lead to the discovery of various spintronic phenomena, including spin pumping—a process in which magnetization dynamics generates a spin current RevModPhys.87.1213 .
While spintronics has primarily focused on the generation and conversion of spin currents, recent advances have revealed that charge currents can also drive the nonequilibrium dynamics of orbital angular momentum (OAM). A key phenomenon underlying OAM dynamics is the direct orbital Hall effect (OHE)—the orbital counterpart of the direct SHE—in which a charge current generates an orbital current, i.e., a flow of electrons carrying nonzero OAM (see Fig. 1)PhysRevLett.95.066601 ; PhysRevLett.102.016601 ; PhysRevLett.121.086602 ; PhysRevB.102.035409 ; PhysRevLett.126.056601 . Theoretical and experimental progress on the OHE has led to the emergence of orbitronics, which aims to utilize the interplay between charge and orbital currents in solid-state devices go2021orbitronics ; KIM2022169974 .
Figure 1: Schematic illustration of orbital torque induced by the direct OHE (left) and the inverse OHE induced by orbital pumping (right) in W/Ni. In the direct OHE, $j_{\mathrm{C}}$ represents the applied charge current, while $j_{\mathrm{L}}^{\mathrm{bulk}}$ and $j_{\mathrm{L}}^{\mathrm{surface}}$ denote the orbital currents generated by the bulk and surface contributions, respectively. The red and blue colors represent the positive and negative orbital polarizations, respectively. In the direct OHE, $j_{\mathrm{L}}^{\mathrm{bulk}}$ and $j_{\mathrm{L}}^{\mathrm{surface}}$ carry opposite orbital angular momentum because the bulk and surface orbital Hall conductivities have opposite signs in W. In the inverse OHE, $j_{\mathrm{L}}$ is the injected orbital current, while $j_{\mathrm{C}}^{\mathrm{bulk}}$ and $j_{\mathrm{C}}^{\mathrm{surface}}$ are the charge currents generated by the bulk and surface contributions, respectively. These charge currents, $j_{\mathrm{C}}^{\mathrm{bulk}}$ and $j_{\mathrm{C}}^{\mathrm{surface}}$, flow in opposite directions. $M$ denotes the magnetization.
[…]
$\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)\neq\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z).$ (1)
This prediction indicates that while the direct and inverse OHEs are globally reciprocal, their local reciprocity is violated. This is in stark contrast to the SHE, which has been predicted to be reciprocal even locally go2024local . Such a distinction is of fundamental importance, as it provides a key experimental signature for differentiating between spin and orbital currents. However, despite recent observations of the direct and inverse OHEs Cr-orbital ; Ta-orbital ; PhysRevResearch.4.033037 ; hayashi2022observation ; choi2021observation ; hayashi2024observation ; hayashi2022observation ; hayashi2023orbital ; choi2021observation ; PhysRevB.107.134423 ; xu2024orbitronics ; hayashi2024observation ; seifert2023time ; gao2024control ; 4c06056 , the nature of their reciprocal relationship has remained unclear. In nanoscale spintronic and orbitronic devices, the measured quantities do not exactly correspond to macroscopic averages. In such devices, local reciprocity may be investigated by carefully measuring spin and orbital responses across devices with systematically varied thicknesses. This approach could offer a pathway to exploring coupled transport phenomena beyond the framework of Onsager’s reciprocity.
In this Letter, we report the observation of a violation of local reciprocity in the charge-orbital interconversion. We investigate the direct and inverse OHEs in W using orbital torque and orbital pumping in W/Ni bilayers (see Fig. 1PhysRevResearch.2.013177 ; hayashi2024observation . In this system, the charge-to-spin/orbital conversion can be investigated by measuring current-induced spin and orbital torques on the Ni magnetization hayashi2022observation . The reciprocal process, the spin/orbital-to-charge conversion is induced by ferromagnetic resonance (FMR). The FMR drives the orbital pumping, as well as the spin pumping, due to the strong spin-orbit correlation of Ni hayashi2024observation ; arXiv:2309.14817 ; han2023theory ; PhysRevApplied.19.014069 ; el2023observation , injecting spin and orbital currents into the W layer. These currents are subsequently converted into charge currents via the inverse SHE and OHE, respectively. Through systematic measurements of the orbital torque and orbital pumping, we demonstrate that the OHE in the W bulk exhibits opposite signs in the direct and inverse processes. This asymmetry aligns with a theoretical prediction of the strong local non-reciprocity in the OHEs of W go2024local , revealing the unique characteristics of charge-orbital coupled transport.
To investigate the charge-to-orbital conversion in W, we measured the current-induced torque in the W/Ni bilayers using spin-torque ferromagnetic resonance (ST-FMR). The samples were fabricated by magnetron sputtering with the structure of SiO2(4 nm)/W($t_{\mathrm{W}}$)/Ni($t_{\mathrm{Ni}}$)/SiO2-substrate , where the numbers in parentheses indicate the layer thicknesses. The measurement and analysis procedures for the ST-FMR follow those in our previous work hayashi2022observation . In Fig. 2(a), we show the W-layer thickness $t_{\mathrm{W}}$ dependence of the damping-like torque efficiency, $\xi_{\text{DL}}^{E}=(2e/\hbar)\mu_{0}M_{\text{s}}t_{\text{Ni}}H_{\text{DL}}/E$, determined by the ST-FMR for the W/Ni bilayers, where $M_{\mathrm{s}}$ is the saturation magnetization, $H_{\mathrm{DL}}$ is the damping-like effective field, and $E$ is the applied electric field.
Figure 2: The damping-like torque efficiency $\xi_{\mathrm{DL}}^{E}$ determined by the ST-FMR for the W/Ni bilayers. (a) Dependence of $\xi_{\mathrm{DL}}^{E}$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer. (b) Dependence of $\xi_{\mathrm{DL}}^{E}$ on the Ni thickness $t_{\rm Ni}$ in the W(3 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers.
Figure 2(a) shows that $\xi_{\text{DL}}^{E}$ changes sign from negative to positive as $t_{\mathrm{W}}$ increases, indicating a crossover in the dominant mechanism responsible for the observed torque. This result is consistent with our previous observation hayashi2022observation . The negative sign of $\xi_{\mathrm{DL}}^{E}$ aligns with that of the SHE in the W layer PhysRevMaterials.6.095001 , suggesting the direct SHE is the primary mechanism driving this torque. We note that the increase in $\xi_{\mathrm{DL}}^{E}$ with $t_{\mathrm{W}}$ indicates that the torque with a positive sign becomes pronounced as $t_{\mathrm{W}}$ increases, suggesting that this torque originates from the charge current flowing in the bulk of the W layer. This $t_{\mathrm{W}}$-dependent behavior is consistent with what is expected from the bulk direct OHE in the W layer.
The significant contribution of the orbital torque to $\xi_{\mathrm{DL}}^{E}$ in the W/Ni bilayer is supported by a long-range nature of $\xi_{\mathrm{DL}}^{E}$. A key difference between spin and orbital phenomena lies in the characteristic length scales associated with the thickness of the ferromagnet (FM) PhysRevLett.130.246701 . It has been demonstrated that spin torque efficiency is nearly independent of the FM-layer thickness due to the short-range nature of spin transport in FMs, whereas orbital torque efficiency increases with FM-layer thickness because of its long-range nature PhysRevLett.130.246701 ; hayashi2022observation ; PhysRevLett.128.067201 ; PhysRevB.107.134423 ; PhysRevResearch.4.033037 ; PhysRevB.105.104434 ; PhysRevResearch.5.023054 ; gao2024control . This difference in the length scales arises because spin currents decay rapidly—typically within less than 1 nm–due to spin dephasing in FMs, whereas orbital currents can propagate over much longer distances through momentum-space hotspots, which suppress oscillations of the OAM PhysRevLett.130.246701 . As shown in Fig. 2(b), $\xi_{\mathrm{DL}}^{E}$ increases with $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer, even though the $t_{\mathrm{Ni}}$ range is much larger than the spin dephasing length. This result suggests that $\xi_{\mathrm{DL}}^{E}$ with the positive sign is dominated by the orbital torque induced by the bulk direct OHE. Here, Figure 2(b) also shows that in the W(3 nm)/Ni($t_{\mathrm{Ni}}$) bilayer, $\xi_{\mathrm{DL}}^{E}$ increases with $t_{\mathrm{Ni}}$ and changes sign. This result can be interpreted as arising from a competition between spin and orbital torques. While the spin torque due to the direct SHE with the negative sign provides the dominant contribution to $\xi_{\mathrm{DL}}^{E}$ in the W(3 nm)/Ni(5 nm) bilayer, this contribution remains nearly independent of $t_{\mathrm{Ni}}$. In contrast, the orbital torque with the positive sign due to the direct OHE increases with $t_{\mathrm{Ni}}$, leading to the sign reversal of $\xi_{\mathrm{DL}}^{E}$.
The W layer in the W/Ni bilayer predominantly consists of $\alpha$-W, which is confirmed by the dependence of the W-layer resistivity $\rho_{\rm W}$ on $t_{\rm W}$, shown in Fig. 3(a). Figure 3(a) shows that $\rho_{\rm W}$ follows $\rho_{\rm W}(t_{\rm W})=at_{\rm W}^{-1}+\rho_{\rm W}^{\rm bulk}$, where $at_{\rm W}^{-1}$ accounts for the resistivity contribution from surface scattering, and $\rho_{\rm W}^{\rm bulk}$ is the bulk-limit resistivity. The extracted value of $\rho_{\rm W}^{\rm bulk}=17~{}\mu\Omega\text{cm}$ is consistent with the resistivity of low-resistivity $\alpha$-W hayashi2022observation . These results indicate that the sign of the bulk direct orbital Hall conductivity $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}$ is positive in $\alpha$-W. This observation is consistent with theoretical predictions PhysRevLett.102.016601 ; PhysRevMaterials.6.095001 ; go2024local .
In the following, we explore the orbital-to-charge conversion in the $\alpha$-W layer by investigating the inverse process of the orbital torque generation—orbital pumping. For the pumping measurements, the W/Ni film was patterned into rectangular strips with a width of $8$ $\mu$m and a length of $180$ $\mu$m using photolithography and Ar-ion milling. The device was positioned between the ground and signal lines of a Ti/Au coplanar waveguide (CPW), with a gap of 8 $\mu$m separating the CPW from the device (see Fig. 3(b)). During the measurements, a radio frequency (RF) current $I_{\mathrm{RF}}$ with a frequency $f$ and a power $P_{\mathrm{in}}$ was applied to the CPW, along with an in-plane external magnetic field $H$ at an angle $\theta$ relative to the longitudinal direction of the device. The RF current generates an out-of-plane microwave magnetic field, driving FMR in the W/Ni bilayer. We measured the direct-current (DC) voltage $V_{\rm DC}$ induced by the FMR across electrodes attached to the W/Ni bilayer.
Figure 3: Charge-current signals induced by spin/orbital pumping. (a) Dependence of the W-layer resistivity $\rho_{\mathrm{W}}$ on $t_{\rm W}$ in the Ni(5 nm)/W($t_{\mathrm{W}}$) bilayer. The solid circles are the experimental data, and the solid cure is the fitting result. (b) Schematic illustration of the experimental setup used to measure the spin/orbital pumping. ${\bf H}$ and ${\bf I}_{\mathrm{RF}}$ denote the external magnetic field and the RF current, respectively. $\theta$ is defined as the in-plane angle between ${\bf H}$ and ${\bf I}_{\mathrm{RF}}$. $H$ dependence of $I_{\rm DC}$ for the (c) W(0.4 nm)/Ni(5 nm) and (d) W(6 nm)/Ni(5 nm) bilayers at $\theta=90^{\circ}$, with $f=10$ GHz and $P_{\rm in}=200$ mW. The solid circles represent the experimental data, and the black solid curves show the fitting results, which are composed of the sum of symmetric (red curve) and antisymmetric (blue curve) components. $\theta$ dependence of $I_{\rm{sym}}$ for the (e) W(0.4 nm)/Ni(5 nm) and (f) W(6 nm)/Ni(5 nm) bilayers. The solid circles represent the experimental data, and the black solid curves show the fitting results using Eq. (2).
In Figs. 3(c) and 3(d), we show the dependence of the charge current $I_{\rm DC}=V_{\rm DC}/R$ on $H$ for the W(0.4 nm)/Ni and W(6 nm)/Ni bilayers, respectively, where $R$ denotes the device resistance. By fitting the measured signal using $I_{\rm{DC}}=I_{\rm{sym}}{(\mu_{0}W)^{2}}/[{(\mu_{0}H-\mu_{0}H_{\rm{res}})^{2}+(\mu_{0}W)^{2}}]+I_{\rm{asym}}{(\mu_{0}W)(\mu_{0}H-\mu_{0}H_{\rm{res}})}/[{(\mu_{0}H-\mu_{0}H_{\rm{res}})^{2}+(\mu_{0}W)^{2}}]$, we extract the symmetric component $I_{\rm{sym}}$, where $H_{\rm{res}}$ is the FMR field and $W$ is the FMR linewidth. Here, the spin and orbital pumping generates $I_{\rm{sym}}$, while spin rectification effects, including the anisotropic magnetoresistance (AMR), the planar Hall effect (PHE), and the anomalous Hall effect (AHE) in the Ni layer, can contribute to both the symmetric $I_{\rm{sym}}$ and antisymmetric $I_{\rm{asym}}$ components.
To extract the pumping signal $I_{\mathrm{pump}}$, we measured $I_{\mathrm{DC}}$ by varying the in-plane magnetic field angle $\theta$, as shown in Figs. 3(e) and 3(f). The results are consistent with HARDER20161
$I_{\mathrm{sym}}=I_{\mathrm{pump}}\sin\theta+I_{\mathrm{AMR}}\sin 2\theta+I_{\mathrm{PHE}}\cos 2\theta+I_{\mathrm{AHE}},$ (2)
where $I_{\mathrm{AHR}}$, $I_{\mathrm{PHE}}$, and $I_{\mathrm{AHE}}$ are the spin rectification signals due to the AMR, PHE, and AHE, respectively. Figures 3(e) and 3(f) demonstrate that the sizable $I_{\mathrm{pump}}$ signals are generated by the spin and/or orbital pumping in the W/Ni bilayers. Using the extracted value of $I_{\mathrm{pump}}$, we determine the pumping-induced charge-current generation efficiency $\zeta$, defined as supplementary
$\zeta=I_{\mathrm{pump}}\frac{\gamma\mu_{0}H_{\mathrm{res}}\left({\mu_{0}W}\right)^{2}}{2\pi ewf^{2}}\left[1+\left(\frac{2\pi f}{\gamma\mu_{0}H_{\mathrm{res}}}\right)^{2}\right]^{2},$ (3)
where $\gamma$ is the gyromagnetic ratio, $e$ is the elementary charge, and $w$ is the width of the device. In the model of the spin pumping and the inverse SHE, $\zeta$ represents an efficiency parameter that characterizes the strength of the spin pumping and the inverse SHE as $\zeta=g_{\mathrm{eff}}^{\uparrow\downarrow}\theta_{\mathrm{H}}\lambda\tanh\left[{t_{\mathrm{W}}}/(2\lambda)\right](\mu_{0}h)^{2}$, which varies with the W thickness over the spin diffusion length but should be independent of the Ni thickness. Here, $g_{\mathrm{eff}}^{\uparrow\downarrow}$ is the effective spin-mixing conductance, $\theta_{\mathrm{H}}$ is the spin Hall angle, and $\lambda$ is the spin diffusion length. $h$ is the amplitude of the microwave magnetic field, which is consistent across all devices used in this study.
Figure 4: The pumping-induced charge-current generation efficiency $\zeta$ in the W/Ni bilayers. (a) Dependence of $\zeta$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer. (b) Dependence of $\zeta$ on the Ni thickness $t_{\rm Ni}$ in the W(0.6 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers.
In Fig. 4(a), we show the dependence of the pumping efficiency $\zeta$ on the W-layer thickness $t_{\mathrm{W}}$. Notably, $\zeta$ remains negative in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer even for $t_{\mathrm{W}}>6$ nm, where the orbital torque efficiency $\xi_{\mathrm{DL}}^{E}$ becomes positive (see Fig. 2(a)). The negative sign of $\zeta$ corresponds to the negative sign of the charge-current generation mechanism in the W layer. This implies that $\zeta$ at $t_{\mathrm{W}}>6$ nm is dominated by the inverse SHE, which is known to be negative in W. Nevertheless, as shown in Fig. 4(b), $\zeta$ exhibits a strong dependence on the Ni-layer thickness $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer. This behavior demonstrates that $\zeta$ cannot be solely attributed to the spin pumping, which should be independent of $t_{\mathrm{Ni}}$, just as the spin torque—the reciprocal process of the spin pumping—is also independent of $t_{\mathrm{Ni}}$. We further confirm that the pumping signal is negligible in Ni single-layer films within this $t_{\mathrm{Ni}}$ range supplementary .
The observed long-range dependence of $\zeta$ on $t_{\mathrm{Ni}}$ is characteristic of the reciprocal process of the long-range orbital torque, indicating that the orbital pumping contributes to the charge current in the W(10 nm)/Ni bilayer. Notably, the decrease in $\zeta$ with $t_{\mathrm{Ni}}$ indicates that the orbital-to-charge conversion has a negative sign. In addition to this negative orbital-to-charge conversion, we also observe a positive orbital-to-charge conversion in the W/Ni bilayer. As shown in Fig. 4(a), $\zeta$ exhibits a positive peak around $t_{\mathrm{W}}=0.5$ nm. This positive component increases with $t_{\mathrm{Ni}}$ as shown in Fig. 4(b), indicating that the signal arises from the orbital-to-charge conversion with a positive sign induced by the orbital pumping. These results demonstrate that two distinct orbital-to-charge conversion mechanisms with opposite signs coexist in the W/Ni bilayer: a positive contribution for $t_{\mathrm{W}}<1$ nm and a negative contribution for $t_{\mathrm{W}}>1$ nm.
The orbital-to-charge conversion with the negative sign can be attributed to the bulk inverse OHE in the W layer, demonstrating a violation of local reciprocity between the direct and inverse OHEs. The nature of the orbital-to-charge conversion can differ between the bulk and interfaces, accounting for the two orbital-to-charge conversion mechanisms with different signs observed in the W/Ni bilayer. Since the interface contribution originates from the anisotropic crystal potential, which differs from the bulk crystal potential go2024local , we assume that the interface contribution is particularly significant at the SiO2/W interface rather than at the metallic W/Ni interface. This interface effect should be especially strong at small $t_{\mathrm{W}}$, suggesting that the charge-current generation with the positive sign at $t_{\mathrm{W}}<1$ nm can be attributed to the orbital-to-charge conversion at the SiO2/W interface. As $t_{\mathrm{W}}$ increases, this surface contribution is suppressed, while bulk effects, including the bulk inverse SHE and OHE, become more pronounced. This indicates that the bulk inverse OHE dominates the observed orbital-to-charge conversion with the negative sign at larger $t_{\mathrm{W}}$. These results demonstrate that while the sign of the bulk direct OHE is positive, as evidenced by the torque measurements, the pumping measurements show that the sign of the bulk inverse OHE is negative. The observation of opposite signs for the bulk direct and inverse OHEs demonstrates that the local reciprocity is violated in the charge-orbital interconversion in the $\alpha$-W.
Table 1: Decomposition of the bulk and surface contributions to the direct and inverse OHEs in W go2024local . The bulk and surface contributions to the direct(inverse) orbital Hall conductivity are represented as $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),surface}}$, respectively, along with the spin Hall conductivity components represented as $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),surface}}$.
Orbital/Spin Conversion Bulk Surface
OHE Direct $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}<0$
Inverse $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{OH}}^{\mathrm{inv,surface}}>0$
SHE Direct $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}\simeq 0$
Inverse $\sigma_{\mathrm{SH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{inv,surface}}\simeq 0$
Our observations are consistent with the theoretical prediction of a violation of local reciprocity between the direct and inverse OHEs in $\alpha$-W go2024local . In the theory, the non-reciprocity in the charge-orbital interconversion arises from the non-conservation of the OAM due to its strong interaction with the lattice via crystal fields. This interaction is accounted for by defining the orbital current as the sum of the conventional orbital current and the torque dipole, the latter being associated with the interaction between the OAM and the crystal lattice. It has been shown that the torque dipole is particularly significant at interfaces, where the anisotropic crystal potential efficiently mediates angular momentum exchange between the electrons and the lattice, strongly influencing the local profile of the charge-orbital interconversion in the W layer. As a result, the local charge currents induced by the orbital current deviate from the expectation of strict reciprocity. Notably, even the signs of the local responses differ between the direct and inverse OHEs, as summarized in Table I. The positive and negative signs of the bulk direct and inverse orbital Hall conductivities, $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ and $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$, evidenced by the torque and pumping results, are consistent with this prediction. Importantly, these opposite signs do not violate Onsager’s reciprocity relation, as the direct and inverse responses are reciprocal when summed over the entire film thickness go2024local . These behaviors contrasts sharply with the SHE in $\alpha$-W, where the direct and inverse SHEs are largely reciprocal even at a local level go2024local . This indicates that the observed violation of the local reciprocity between the direct and inverse OHEs highlights the distinct nature of the charge-orbital interconversion compared to the charge-spin interconversion. Table I also suggests that the surface direct OHE with a negative sign may contribute to the negative torque for $t_{\mathrm{W}}<4$ nm.
In summary, we have experimentally demonstrated the violation of local reciprocity in the charge-orbital interconversion in the W/Ni bilayers. By measuring the orbital torque and orbital pumping, we found that the sign of the bulk inverse OHE is opposite to that of the bulk direct OHE. This result confirms that the charge-orbital interconversion is locally non-reciprocal, in sharp contrast to the global reciprocity governed by Onsager’s reciprocal relations. The observation of the local reciprocity violation in the direct and inverse OHEs highlights the unique characteristics of charge-orbital coupled transport, providing essential insights for a deeper understanding of orbitronic phenomena.
Acknowledgements.
This work was supported by JSPS KAKENHI (Grant Number: 22H04964, 23K19037), Spintronics Research Network of Japan (Spin-RNJ), and MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) (Grant Number: JPJ011438). Y.M. and D.G. acknowledge support from the EIC Pathfinder OPEN grant 101129641 “OBELIX”.

### v2

@@
 By investigating orbital torque and orbital pumping in W/Ni bilayers, we show that the charge-orbital interconversion in the bulk of the W layer exhibits opposite signs in the direct and inverse processes—the direct and inverse orbital Hall effects being positive and negative, respectively.
-This finding provides direct evidence of local non-reciprocity in the charge-orbital interconversion, in agreement with a theoretical prediction.
+This finding provides direct evidence of local nonreciprocity in the charge-orbital interconversion, in agreement with a theoretical prediction.
 These results highlight the unique characteristics of charge-orbital coupled transport and offer fundamental insights into the mechanisms underlying orbital-current-driven phenomena.
@@
 The direct process of the SHE converts a charge current into a spin current, enabling the electric control of magnetization through spin-orbit torques AndoPRL ; liu2012spin .
-This process has attracted extensive attention for its topological and quantum mechanical nature, as well as its potential for spintronic applications, such as non-volatile memories and neuromorphic computing RevModPhys.91.035004 .
+This process has attracted extensive attention for its topological and quantum mechanical nature, as well as its potential for spintronic applications RevModPhys.91.035004 .
 The inverse process of the SHE, the conversion from a spin current into a charge current, offers a way for the electric detection of spin currents saitoh2006conversion ; valenzuela2006direct ; KimuraPRL .
-This mechanism has lead to the discovery of various spintronic phenomena, including spin pumping—a process in which magnetization dynamics generates a spin current RevModPhys.87.1213 .
+This mechanism has led to the discovery of various spintronic phenomena, including spin pumping—a process in which magnetization dynamics generates a spin current RevModPhys.87.1213 .
 While spintronics has primarily focused on the generation and conversion of spin currents, recent advances have revealed that charge currents can also drive the nonequilibrium dynamics of orbital angular momentum (OAM).
 A key phenomenon underlying OAM dynamics is the direct orbital Hall effect (OHE)—the orbital counterpart of the direct SHE—in which a charge current generates an orbital current, i.e., a flow of electrons carrying nonzero OAM (see Fig.
-1)PhysRevLett.95.066601 ; PhysRevLett.102.016601 ; PhysRevLett.121.086602 ; PhysRevB.102.035409 ; PhysRevLett.126.056601 .
-Theoretical and experimental progress on the OHE has led to the emergence of orbitronics, which aims to utilize the interplay between charge and orbital currents in solid-state devices go2021orbitronics ; KIM2022169974 .
+1PhysRevLett.95.066601 ; PhysRevLett.102.016601 ; PhysRevLett.121.086602 ; PhysRevB.102.035409 ; PhysRevLett.126.056601 .
+Theoretical and experimental progress on the direct and inverse OHEs has led to the emergence of orbitronics, which aims to utilize the interplay between charge and orbital currents in solid-state devices go2021orbitronics ; KIM2022169974 ; 2025Andoreview .
 Figure 1: Schematic illustration of orbital torque induced by the direct OHE (left) and the inverse OHE induced by orbital pumping (right) in W/Ni.
@@
 $M$ denotes the magnetization.
-[…]
-$\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)\neq\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z).$ (1)
-This prediction indicates that while the direct and inverse OHEs are globally reciprocal, their local reciprocity is violated.
-This is in stark contrast to the SHE, which has been predicted to be reciprocal even locally go2024local .
-Such a distinction is of fundamental importance, as it provides a key experimental signature for differentiating between spin and orbital currents.
-However, despite recent observations of the direct and inverse OHEs Cr-orbital ; Ta-orbital ; PhysRevResearch.4.033037 ; hayashi2022observation ; choi2021observation ; hayashi2024observation ; hayashi2022observation ; hayashi2023orbital ; choi2021observation ; PhysRevB.107.134423 ; xu2024orbitronics ; hayashi2024observation ; seifert2023time ; gao2024control ; 4c06056 , the nature of their reciprocal relationship has remained unclear.
-In nanoscale spintronic and orbitronic devices, the measured quantities do not exactly correspond to macroscopic averages.
-In such devices, local reciprocity may be investigated by carefully measuring spin and orbital responses across devices with systematically varied thicknesses.
-This approach could offer a pathway to exploring coupled transport phenomena beyond the framework of Onsager’s reciprocity.
-In this Letter, we report the observation of a violation of local reciprocity in the charge-orbital interconversion.
-We investigate the direct and inverse OHEs in W using orbital torque and orbital pumping in W/Ni bilayers (see Fig.
-1PhysRevResearch.2.013177 ; hayashi2024observation .
-In this system, the charge-to-spin/orbital conversion can be investigated by measuring current-induced spin and orbital torques on the Ni magnetization hayashi2022observation .
-The reciprocal process, the spin/orbital-to-charge conversion is induced by ferromagnetic resonance (FMR).
-The FMR drives the orbital pumping, as well as the spin pumping, due to the strong spin-orbit correlation of Ni hayashi2024observation ; arXiv:2309.14817 ; han2023theory ; PhysRevApplied.19.014069 ; el2023observation , injecting spin and orbital currents into the W layer.
-These currents are subsequently converted into charge currents via the inverse SHE and OHE, respectively.
-Through systematic measurements of the orbital torque and orbital pumping, we demonstrate that the OHE in the W bulk exhibits opposite signs in the direct and inverse processes.
-This asymmetry aligns with a theoretical prediction of the strong local non-reciprocity in the OHEs of W go2024local , revealing the unique characteristics of charge-orbital coupled transport.
-To investigate the charge-to-orbital conversion in W, we measured the current-induced torque in the W/Ni bilayers using spin-torque ferromagnetic resonance (ST-FMR).
+One of the fundamental challenges in harnessing charge-orbital coupled transport is to understand the reciprocity of charge-orbital interconversion PhysRevLett.132.226704 ; 4c06056 ; gao2025nonlocalelectricaldetectionreciprocal .
+In the OHE, the local, layer-resolved charge-to-orbital and orbital-to-charge conversion processes are characterized by the local direct and inverse orbital Hall conductivities, $\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)$ and $\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z)$, at position $z$, with the orbital current flowing along the $z$ direction supplementary .
+The global response, obtained by integrating over the film thickness, is governed by Onsager’s reciprocal relations, which require
+$\int\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)dz=\int\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z)dz,$ (1)
+a relation that has been rigorously proven for the proper orbital current go2024local .
+On the other hand, local reciprocity, which concerns whether the equality $\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)=\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z)$ holds at each position $z$, is nontrivial.
+This is because Onsager’s reciprocal relations are derived as a corollary of the fluctuation-dissipation theorem, which relates macroscopic responses in nonequilibrium to microscopic correlations in equilibrium.
+While this guarantees reciprocity for global, spatially integrated responses, it does not impose constraints on local responses.
+In fact, first-principles calculations for W have explicitly demonstrated a violation of local reciprocity between direct and inverse processes,
+$\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)\neq\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z),$ (2)
+even though the global reciprocity condition in Eq.
+(1) is satisfied go2024local .
+This local nonreciprocity originates from the non-conservation of OAM due to its strong coupling with the lattice through crystal fields.
+Such coupling is accounted for by defining the orbital current as the sum of the conventional orbital current and torque dipole, the latter being associated with the interaction between OAM and the crystal lattice go2024local .
+The torque dipole is particularly significant at interfaces, where the anisotropic crystal potential efficiently mediates angular momentum exchange between the electrons and the lattice, strongly influencing the local profile of charge-orbital interconversion in the W layer.
+As a consequence, local charge currents induced by orbital currents deviate from the expectation of strict reciprocity.
+Notably, in $\alpha$-W, the local responses exhibit opposite signs between the direct and inverse OHEs, as summarized in Table I in Appendix A.
+However, despite recent observations of the direct and inverse OHEs Cr-orbital ; Ta-orbital ; PhysRevResearch.4.033037 ; hayashi2022observation ; choi2021observation ; hayashi2022observation ; hayashi2023orbital ; PhysRevB.107.134423 ; xu2024orbitronics ; hayashi2024observation ; seifert2023time ; gao2024control ; 4c06056 , the nature of their reciprocal relationship has remained unclear.
+In this Letter, we report the observation of a violation of local reciprocity in the charge-orbital interconversion in $\alpha$-W.
+Because of the predicted opposite signs between the bulk direct and inverse orbital Hall conductivities go2024local , W is a suitable platform for testing local nonreciprocity supplementary .
+In nanoscale spintronic and orbitronic devices, where measured quantities can deviate from macroscopic averages, local reciprocity can be investigated by systematically varying the device thickness.
+This approach provides a route to exploring coupled transport phenomena beyond the framework of Onsager’s reciprocity.
+Through systematic measurements of orbital torque and orbital pumping in W/Ni bilayers, we demonstrate that the OHE in the W bulk exhibits opposite signs in the direct and inverse processes, revealing the unique nature of charge-orbital coupled transport.
+We measured the current-induced torque in W/Ni bilayers using spin-torque ferromagnetic resonance (ST-FMR) to investigate the charge-to-orbital conversion in W.
 The samples were fabricated by magnetron sputtering with the structure of SiO2(4 nm)/W($t_{\mathrm{W}}$)/Ni($t_{\mathrm{Ni}}$)/SiO2-substrate, where the numbers in parentheses indicate the layer thicknesses.
-The measurement and analysis procedures for the ST-FMR follow those in our previous work hayashi2022observation .
-In Fig.
-2(a), we show the W-layer thickness $t_{\mathrm{W}}$ dependence of the damping-like torque efficiency, $\xi_{\text{DL}}^{E}=(2e/\hbar)\mu_{0}M_{\text{s}}t_{\text{Ni}}H_{\text{DL}}/E$, determined by the ST-FMR for the W/Ni bilayers, where $M_{\mathrm{s}}$ is the saturation magnetization, $H_{\mathrm{DL}}$ is the damping-like effective field, and $E$ is the applied electric field.
-Figure 2: The damping-like torque efficiency $\xi_{\mathrm{DL}}^{E}$ determined by the ST-FMR for the W/Ni bilayers.
+The measurement and analysis procedures follow those in our previous work supplementary ; hayashi2022observation .
+Figure 2(a) shows the W-layer thickness $t_{\mathrm{W}}$ dependence of the damping-like torque efficiency,
+$\xi_{\text{DL}}^{E}=\left(\frac{2e}{\hbar}\right)\frac{\mu_{0}M_{\text{s}}t_{\text{Ni}}H_{\text{DL}}}{E},$ (3)
+determined by ST-FMR for the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayers, where $e$ is the elementary charge, $\hbar$ is the reduced Planck constant, $M_{\mathrm{s}}$ is the saturation magnetization, $H_{\mathrm{DL}}$ is the damping-like effective field, and $E$ is the applied electric field.
+At $t_{\mathrm{W}}<5$ nm, $\xi_{\text{DL}}^{E}$ with the negative sign appears as $t_{\mathrm{W}}$ increases.
+This behavior is consistent with the torque arising from the direct SHE in the W layer.
+In a simple spin diffusion model, the SHE-induced $\xi_{\text{DL}}^{E}$ scales as $1-\mathrm{sech}(t_{\mathrm{W}}/\lambda_{\mathrm{S}})$, where $\lambda_{\mathrm{S}}$ is the spin diffusion length PhysRevLett.106.036601 .
+This model predicts that SHE-induced $\xi_{\text{DL}}^{E}$ increases with $t_{\mathrm{W}}$ and saturates when $t_{\mathrm{W}}$ exceeds approximately $3\lambda_{\mathrm{S}}$.
+The negative sign of $\xi_{\text{DL}}^{E}$ aligns with that of the direct SHE in the W layer PhysRevMaterials.6.095001 , and the rapid increase in $\xi_{\text{DL}}^{E}$ with $t_{\mathrm{W}}$ is consistent with the short spin diffusion length of W PhysRevMaterials.2.014403 , suggesting that the direct SHE is the primary mechanism driving the torque at $t_{\mathrm{W}}<5$ nm.
+Figure 2: The damping-like torque efficiency $\xi_{\mathrm{DL}}^{E}$ determined by ST-FMR for the W/Ni bilayers.
 (a) Dependence of $\xi_{\mathrm{DL}}^{E}$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer.
+The values of $t_{\mathrm{W}}$, determined from the sputtering rate, are as follows: 0, 0.61, and 0.79 nm; in 0.36 nm steps from 1.15 to 2.24 nm; in 0.73 nm steps from 2.97 to 8.06 nm; and 9.15 nm.
 (b) Dependence of $\xi_{\mathrm{DL}}^{E}$ on the Ni thickness $t_{\rm Ni}$ in the W(3 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers.
+The values of $t_{\mathrm{Ni}}$, determined from the sputtering rate, range from 4.81 to 10.58 nm in steps of 1.92 nm.
 Figure 2(a) shows that $\xi_{\text{DL}}^{E}$ changes sign from negative to positive as $t_{\mathrm{W}}$ increases, indicating a crossover in the dominant mechanism responsible for the observed torque.
-This result is consistent with our previous observation hayashi2022observation .
-The negative sign of $\xi_{\mathrm{DL}}^{E}$ aligns with that of the SHE in the W layer PhysRevMaterials.6.095001 , suggesting the direct SHE is the primary mechanism driving this torque.
-We note that the increase in $\xi_{\mathrm{DL}}^{E}$ with $t_{\mathrm{W}}$ indicates that the torque with a positive sign becomes pronounced as $t_{\mathrm{W}}$ increases, suggesting that this torque originates from the charge current flowing in the bulk of the W layer.
-This $t_{\mathrm{W}}$-dependent behavior is consistent with what is expected from the bulk direct OHE in the W layer.
-The significant contribution of the orbital torque to $\xi_{\mathrm{DL}}^{E}$ in the W/Ni bilayer is supported by a long-range nature of $\xi_{\mathrm{DL}}^{E}$.
+Complementary second-harmonic Hall measurements confirm the reliability of $\xi_{\mathrm{DL}}^{E}$ determined by ST-FMR, including the sign reversal with increasing $t_{\mathrm{W}}$ hayashi2022observation .
+At $t_{\mathrm{W}}>5$ nm, $\xi_{\text{DL}}^{E}$ increases with $t_{\mathrm{W}}$, suggesting that an additional torque with the positive sign becomes dominant at the larger $t_{\mathrm{W}}$.
+The observed $t_{\mathrm{W}}$-dependent behavior of this additional torque suggests that it originates from the charge current flowing in the bulk of the W layer, rather than from the SiO2/W, W/Ni, and Ni/SiO2 interfaces supplementary .
+We note that the positive sign of the additional torque is consistent with the sign of the direct OHE in W PhysRevMaterials.6.095001 , suggesting that it can be attributed to the bulk direct OHE.
+In analogy to the model of the SHE-induced torque, the orbital torque due to the direct OHE in the W layer may also scale as $1-\mathrm{sech}(t_{\mathrm{W}}/\lambda_{\mathrm{L}})$, where $\lambda_{\mathrm{L}}$ is the characteristic orbital transport length.
+The observed $t_{\mathrm{W}}$-dependence of the additional torque suggests that $\lambda_{\mathrm{L}}$ is larger than $\lambda_{\mathrm{S}}$, consistent with previous reports hayashi2022observation ; PhysRevResearch.4.033037 ; choi2021observation .
+In this case, the orbital Hall contribution to $\xi_{\text{DL}}^{E}$ from the bulk charge current is negligible at small $t_{\mathrm{W}}$.
+However, this contribution increases with $t_{\mathrm{W}}$ and can exceed that from the SHE at sufficiently large $t_{\mathrm{W}}$, resulting in the sign reversal of $\xi_{\text{DL}}^{E}$ with $t_{\mathrm{W}}$.
 A key difference between spin and orbital phenomena lies in the characteristic length scales associated with the thickness of the ferromagnet (FM) PhysRevLett.130.246701 .
-It has been demonstrated that spin torque efficiency is nearly independent of the FM-layer thickness due to the short-range nature of spin transport in FMs, whereas orbital torque efficiency increases with FM-layer thickness because of its long-range nature PhysRevLett.130.246701 ; hayashi2022observation ; PhysRevLett.128.067201 ; PhysRevB.107.134423 ; PhysRevResearch.4.033037 ; PhysRevB.105.104434 ; PhysRevResearch.5.023054 ; gao2024control .
-This difference in the length scales arises because spin currents decay rapidly—typically within less than 1 nm–due to spin dephasing in FMs, whereas orbital currents can propagate over much longer distances through momentum-space hotspots, which suppress oscillations of the OAM PhysRevLett.130.246701 .
+When a spin current is injected into a FM, it decays rapidly—typically within 1 nm—due to spin dephasing.
+Because of this short-range nature of spin transport, the damping-like effective field due to spin injection scales as $H_{\mathrm{DL}}\propto 1/t_{\mathrm{Ni}}$.
+Consequently, the damping-like torque efficiency defined in Eq.
+(3) is independent of the FM-layer thickness once it exceeds the spin dephasing length, as demonstrated in spin-Hall-dominated Pt/Ni81Fe19 bilayers supplementary .
+In contrast, orbital currents can propagate over much longer distances than spin dephasing length through momentum-space hotspots that suppress oscillations of OAM PhysRevLett.130.246701 .
+Owing to this long-range nature of orbital transport in FMs, $\xi_{\mathrm{DL}}^{E}$ arising from orbital torque increases with the FM-layer thickness PhysRevLett.130.246701 ; hayashi2022observation ; PhysRevLett.128.067201 ; PhysRevB.107.134423 ; PhysRevResearch.4.033037 ; PhysRevB.105.104434 ; PhysRevResearch.5.023054 ; gao2024control .
+The significant orbital contribution to $\xi_{\mathrm{DL}}^{E}$ is supported by a long-range nature of $\xi_{\mathrm{DL}}^{E}$.
 As shown in Fig.
 2(b), $\xi_{\mathrm{DL}}^{E}$ increases with $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer, even though the $t_{\mathrm{Ni}}$ range is much larger than the spin dephasing length.
-This result suggests that $\xi_{\mathrm{DL}}^{E}$ with the positive sign is dominated by the orbital torque induced by the bulk direct OHE.
-Here, Figure 2(b) also shows that in the W(3 nm)/Ni($t_{\mathrm{Ni}}$) bilayer, $\xi_{\mathrm{DL}}^{E}$ increases with $t_{\mathrm{Ni}}$ and changes sign.
-This result can be interpreted as arising from a competition between spin and orbital torques.
+In this $t_{\mathrm{Ni}}$ range, we also confirmed that $\xi_{\mathrm{DL}}^{E}$ is negligible in Ni single layers supplementary .
+These results indicate that $\xi_{\mathrm{DL}}^{E}$ with the positive sign is dominated by the orbital torque induced by the bulk direct OHE.
+Figure 2(b) also shows that in the W(3 nm)/Ni($t_{\mathrm{Ni}}$) bilayer, $\xi_{\mathrm{DL}}^{E}$ increases with $t_{\mathrm{Ni}}$ and changes sign.
+This behavior can be interpreted as arising from a competition between spin and orbital torques.
 While the spin torque due to the direct SHE with the negative sign provides the dominant contribution to $\xi_{\mathrm{DL}}^{E}$ in the W(3 nm)/Ni(5 nm) bilayer, this contribution remains nearly independent of $t_{\mathrm{Ni}}$.
 In contrast, the orbital torque with the positive sign due to the direct OHE increases with $t_{\mathrm{Ni}}$, leading to the sign reversal of $\xi_{\mathrm{DL}}^{E}$.
-The W layer in the W/Ni bilayer predominantly consists of $\alpha$-W, which is confirmed by the dependence of the W-layer resistivity $\rho_{\rm W}$ on $t_{\rm W}$, shown in Fig.
-3(a).
-Figure 3(a) shows that $\rho_{\rm W}$ follows $\rho_{\rm W}(t_{\rm W})=at_{\rm W}^{-1}+\rho_{\rm W}^{\rm bulk}$, where $at_{\rm W}^{-1}$ accounts for the resistivity contribution from surface scattering, and $\rho_{\rm W}^{\rm bulk}$ is the bulk-limit resistivity.
-The extracted value of $\rho_{\rm W}^{\rm bulk}=17~{}\mu\Omega\text{cm}$ is consistent with the resistivity of low-resistivity $\alpha$-W hayashi2022observation .
-These results indicate that the sign of the bulk direct orbital Hall conductivity $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}$ is positive in $\alpha$-W.
-This observation is consistent with theoretical predictions PhysRevLett.102.016601 ; PhysRevMaterials.6.095001 ; go2024local .
-In the following, we explore the orbital-to-charge conversion in the $\alpha$-W layer by investigating the inverse process of the orbital torque generation—orbital pumping.
-For the pumping measurements, the W/Ni film was patterned into rectangular strips with a width of $8$ $\mu$m and a length of $180$ $\mu$m using photolithography and Ar-ion milling.
-The device was positioned between the ground and signal lines of a Ti/Au coplanar waveguide (CPW), with a gap of 8 $\mu$m separating the CPW from the device (see Fig.
-3(b)).
-During the measurements, a radio frequency (RF) current $I_{\mathrm{RF}}$ with a frequency $f$ and a power $P_{\mathrm{in}}$ was applied to the CPW, along with an in-plane external magnetic field $H$ at an angle $\theta$ relative to the longitudinal direction of the device.
-The RF current generates an out-of-plane microwave magnetic field, driving FMR in the W/Ni bilayer.
-We measured the direct-current (DC) voltage $V_{\rm DC}$ induced by the FMR across electrodes attached to the W/Ni bilayer.
+A similar competition between spin and orbital torques has been demonstrated in a W/CoFeB structure, where the dominant mechanism depends on the thicknesses of both the W and CoFeB layers chiba2024comparative .
+The critical thicknesses for the crossover from spin-dominated to orbital-dominated torque are smaller in the W/Ni structure than in the W/CoFeB structure.
+This difference can be attributed to the stronger orbital response of Ni compared to CoFeB Ta-orbital , which facilitates more efficient orbital torque generation and leads to a lower threshold thickness for the onset of orbital-dominated behavior.
+These results support the interpretation that the observed sign change of $\xi_{\mathrm{DL}}^{E}$ arises from the competition between the negative spin torque and the positive orbital torque.
+The W layer in the W/Ni bilayer predominantly consists of $\alpha$-W, which is confirmed by the dependence of the W-layer resistivity $\rho_{\rm W}$ on $t_{\rm W}$ (Appendix B) supplementary .
+These results indicate that the sign of the bulk direct orbital Hall conductivity $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}$ is positive in $\alpha$-W, consistent with theoretical predictions PhysRevLett.102.016601 ; PhysRevMaterials.6.095001 ; go2024local .
 Figure 3: Charge-current signals induced by spin/orbital pumping.
-(a) Dependence of the W-layer resistivity $\rho_{\mathrm{W}}$ on $t_{\rm W}$ in the Ni(5 nm)/W($t_{\mathrm{W}}$) bilayer.
-The solid circles are the experimental data, and the solid cure is the fitting result.
-(b) Schematic illustration of the experimental setup used to measure the spin/orbital pumping.
-${\bf H}$ and ${\bf I}_{\mathrm{RF}}$ denote the external magnetic field and the RF current, respectively.
-$\theta$ is defined as the in-plane angle between ${\bf H}$ and ${\bf I}_{\mathrm{RF}}$.
-$H$ dependence of $I_{\rm DC}$ for the (c) W(0.4 nm)/Ni(5 nm) and (d) W(6 nm)/Ni(5 nm) bilayers at $\theta=90^{\circ}$, with $f=10$ GHz and $P_{\rm in}=200$ mW.
+$H$ dependence of $I_{\rm DC}$ for the (a) W(0.4 nm)/Ni(5 nm) and (b) W(6 nm)/Ni(5 nm) bilayers at $\theta=90^{\circ}$, with $f=10$ GHz and $P_{\rm in}=200$ mW.
 The solid circles represent the experimental data, and the black solid curves show the fitting results, which are composed of the sum of symmetric (red curve) and antisymmetric (blue curve) components.
-$\theta$ dependence of $I_{\rm{sym}}$ for the (e) W(0.4 nm)/Ni(5 nm) and (f) W(6 nm)/Ni(5 nm) bilayers.
+In-plane magnetic field angle $\theta$ dependence of $I_{\rm{sym}}$ for the (c) W(0.4 nm)/Ni(5 nm) and (d) W(6 nm)/Ni(5 nm) bilayers.
 The solid circles represent the experimental data, and the black solid curves show the fitting results using Eq.
 (4).
+In the following, we investigate the orbital-to-charge conversion in W using orbital pumping hayashi2024observation ; arXiv:2309.14817 ; han2023theory ; PhysRevApplied.19.014069 ; el2023observation .
+The experimental setup and measurement procedures are described in Appendix C.
+In the W/Ni bilayer, microwave-driven FMR induces orbital and spin pumping owing to the strong spin-orbit correlation of Ni hayashi2024observation ; arXiv:2309.14817 ; han2023theory ; PhysRevApplied.19.014069 ; el2023observation , injecting orbital and spin currents into the W layer that are converted into charge currents via the inverse OHE and SHE, respectively.
 In Figs.
-3(c) and 3(d), we show the dependence of the charge current $I_{\rm DC}=V_{\rm DC}/R$ on $H$ for the W(0.4 nm)/Ni and W(6 nm)/Ni bilayers, respectively, where $R$ denotes the device resistance.
+3(a) and 3(b), we show the dependence of the charge current $I_{\rm DC}=V_{\rm DC}/R$ on $H$ for the W(0.4 nm)/Ni(5 nm) and W(6 nm)/Ni(5 nm) bilayers, respectively, where $V_{\rm DC}$ is the measured direct-current (DC) voltage, and $R$ denotes the device resistance.
 By fitting the measured signal using $I_{\rm{DC}}=I_{\rm{sym}}{(\mu_{0}W)^{2}}/[{(\mu_{0}H-\mu_{0}H_{\rm{res}})^{2}+(\mu_{0}W)^{2}}]+I_{\rm{asym}}{(\mu_{0}W)(\mu_{0}H-\mu_{0}H_{\rm{res}})}/[{(\mu_{0}H-\mu_{0}H_{\rm{res}})^{2}+(\mu_{0}W)^{2}}]$, we extract the symmetric component $I_{\rm{sym}}$, where $H_{\rm{res}}$ is the FMR field and $W$ is the FMR linewidth.
-Here, the spin and orbital pumping generates $I_{\rm{sym}}$, while spin rectification effects, including the anisotropic magnetoresistance (AMR), the planar Hall effect (PHE), and the anomalous Hall effect (AHE) in the Ni layer, can contribute to both the symmetric $I_{\rm{sym}}$ and antisymmetric $I_{\rm{asym}}$ components.
+Here, spin and orbital pumping generate $I_{\rm{sym}}$, while spin rectification effects, including the anisotropic magnetoresistance (AMR), the planar Hall effect (PHE), and the anomalous Hall effect (AHE) in the Ni layer, can contribute to both the symmetric $I_{\rm{sym}}$ and antisymmetric $I_{\rm{asym}}$ components.
 To extract the pumping signal $I_{\mathrm{pump}}$, we measured $I_{\mathrm{DC}}$ by varying the in-plane magnetic field angle $\theta$, as shown in Figs.
-3(e) and 3(f).
-The results are consistent with HARDER20161
+3(c) and 3(d).
+The results are consistent with HARDER20161 ; supplementary
 $I_{\mathrm{sym}}=I_{\mathrm{pump}}\sin\theta+I_{\mathrm{AMR}}\sin 2\theta+I_{\mathrm{PHE}}\cos 2\theta+I_{\mathrm{AHE}},$ (4)
-where $I_{\mathrm{AHR}}$, $I_{\mathrm{PHE}}$, and $I_{\mathrm{AHE}}$ are the spin rectification signals due to the AMR, PHE, and AHE, respectively.
-Figures 3(e) and 3(f) demonstrate that the sizable $I_{\mathrm{pump}}$ signals are generated by the spin and/or orbital pumping in the W/Ni bilayers.
+where $I_{\mathrm{AMR}}$, $I_{\mathrm{PHE}}$, and $I_{\mathrm{AHE}}$ are the spin rectification signals due to the AMR, PHE, and AHE, respectively.
+Figures 3(c) and 3(d) show that sizable $I_{\mathrm{pump}}$ signals are generated by spin and/or orbital pumping in the W/Ni bilayers.
 Using the extracted value of $I_{\mathrm{pump}}$, we determine the pumping-induced charge-current generation efficiency $\zeta$, defined as supplementary
 $\zeta=I_{\mathrm{pump}}\frac{\gamma\mu_{0}H_{\mathrm{res}}\left({\mu_{0}W}\right)^{2}}{2\pi ewf^{2}}\left[1+\left(\frac{2\pi f}{\gamma\mu_{0}H_{\mathrm{res}}}\right)^{2}\right]^{2},$ (5)
-where $\gamma$ is the gyromagnetic ratio, $e$ is the elementary charge, and $w$ is the width of the device.
-In the model of the spin pumping and the inverse SHE, $\zeta$ represents an efficiency parameter that characterizes the strength of the spin pumping and the inverse SHE as $\zeta=g_{\mathrm{eff}}^{\uparrow\downarrow}\theta_{\mathrm{H}}\lambda\tanh\left[{t_{\mathrm{W}}}/(2\lambda)\right](\mu_{0}h)^{2}$, which varies with the W thickness over the spin diffusion length but should be independent of the Ni thickness.
-Here, $g_{\mathrm{eff}}^{\uparrow\downarrow}$ is the effective spin-mixing conductance, $\theta_{\mathrm{H}}$ is the spin Hall angle, and $\lambda$ is the spin diffusion length.
+where $\gamma$ is the gyromagnetic ratio and $w$ is the width of the device.
+In the model of spin pumping and the inverse SHE, $\zeta$ represents an efficiency parameter that characterizes the combined strength of spin pumping and the inverse SHE as $\zeta=g_{\mathrm{eff}}^{\uparrow\downarrow}\theta_{\mathrm{H}}\lambda_{\mathrm{S}}\tanh\left[{t_{\mathrm{W}}}/(2\lambda_{\mathrm{S}})\right](\mu_{0}h)^{2}$, which varies with the W thickness over the spin diffusion length but should be independent of the Ni thickness.
+Here, $g_{\mathrm{eff}}^{\uparrow\downarrow}$ is the effective spin-mixing conductance, and $\theta_{\mathrm{H}}$ is the spin Hall angle.
 $h$ is the amplitude of the microwave magnetic field, which is consistent across all devices used in this study.
@@
 (a) Dependence of $\zeta$ on the W thickness $t_{\mathrm{W}}$ in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer.
+The values of $t_{\mathrm{W}}$, determined from the sputtering rate, are as follows: 0, 0.42, and 0.79 nm; in 0.36 nm steps from 1.15 to 2.24 nm; in 0.73 nm steps from 2.97 to 8.06 nm; and 9.15 and 9.88 nm.
 (b) Dependence of $\zeta$ on the Ni thickness $t_{\rm Ni}$ in the W(0.6 nm)/Ni($t_{\mathrm{Ni}}$) (blue) and W(10 nm)/Ni($t_{\mathrm{Ni}}$) (red) bilayers.
+The values of $t_{\mathrm{Ni}}$, determined from the sputtering rate, are as follows: for the W(0.6 nm)/Ni($t_{\mathrm{Ni}}$), from 5.00 to 13.30 nm in 2.08 nm steps, and 17.45 nm; for the W(10 nm)/Ni($t_{\mathrm{Ni}}$) structure, 4.68 and 5.72 nm, in 2.08 nm steps from 8.83 to 12.98 nm, and 17.13 nm.
 In Fig.
@@
 The negative sign of $\zeta$ corresponds to the negative sign of the charge-current generation mechanism in the W layer.
-This implies that $\zeta$ at $t_{\mathrm{W}}>6$ nm is dominated by the inverse SHE, which is known to be negative in W.
+This suggests that $\zeta$ at $t_{\mathrm{W}}>6$ nm is dominated by the inverse SHE, which is known to be negative in W.
 Nevertheless, as shown in Fig.
 4(b), $\zeta$ exhibits a strong dependence on the Ni-layer thickness $t_{\mathrm{Ni}}$ in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer.
-This behavior demonstrates that $\zeta$ cannot be solely attributed to the spin pumping, which should be independent of $t_{\mathrm{Ni}}$, just as the spin torque—the reciprocal process of the spin pumping—is also independent of $t_{\mathrm{Ni}}$.
-We further confirm that the pumping signal is negligible in Ni single-layer films within this $t_{\mathrm{Ni}}$ range supplementary .
-The observed long-range dependence of $\zeta$ on $t_{\mathrm{Ni}}$ is characteristic of the reciprocal process of the long-range orbital torque, indicating that the orbital pumping contributes to the charge current in the W(10 nm)/Ni bilayer.
+This behavior indicates that $\zeta$ cannot be solely attributed to spin pumping, which should be independent of $t_{\mathrm{Ni}}$, just as $\xi_{\mathrm{DL}}^{E}$ of spin torque—the reciprocal process of spin pumping—is also independent of $t_{\mathrm{Ni}}$.
+This interpretation is further supported by control experiments on spin-dominated Pt/Ni81Fe19 bilayers, showing that $\zeta$ is independent of the FM thickness supplementary .
+We also confirmed that the pumping signal is negligible in Ni single-layer films within this $t_{\mathrm{Ni}}$ range supplementary .
+The observed long-range dependence of $\zeta$ on $t_{\mathrm{Ni}}$ is characteristic of the reciprocal process of long-range orbital torque, indicating that orbital pumping contributes to the charge current in the W(10 nm)/Ni($t_{\mathrm{Ni}}$) bilayer.
 Notably, the decrease in $\zeta$ with $t_{\mathrm{Ni}}$ indicates that the orbital-to-charge conversion has a negative sign.
-In addition to this negative orbital-to-charge conversion, we also observe a positive orbital-to-charge conversion in the W/Ni bilayer.
+In addition to this negative orbital-to-charge conversion, we also observe a positive orbital-to-charge conversion.
 As shown in Fig.
@@
 This positive component increases with $t_{\mathrm{Ni}}$ as shown in Fig.
-4(b), indicating that the signal arises from the orbital-to-charge conversion with a positive sign induced by the orbital pumping.
-These results demonstrate that two distinct orbital-to-charge conversion mechanisms with opposite signs coexist in the W/Ni bilayer: a positive contribution for $t_{\mathrm{W}}<1$ nm and a negative contribution for $t_{\mathrm{W}}>1$ nm.
-The orbital-to-charge conversion with the negative sign can be attributed to the bulk inverse OHE in the W layer, demonstrating a violation of local reciprocity between the direct and inverse OHEs.
-The nature of the orbital-to-charge conversion can differ between the bulk and interfaces, accounting for the two orbital-to-charge conversion mechanisms with different signs observed in the W/Ni bilayer.
-Since the interface contribution originates from the anisotropic crystal potential, which differs from the bulk crystal potential go2024local , we assume that the interface contribution is particularly significant at the SiO2/W interface rather than at the metallic W/Ni interface.
+4(b), indicating that the signal arises from the orbital-to-charge conversion with the positive sign induced by orbital pumping.
+These results demonstrate that two distinct orbital-to-charge conversion mechanisms with opposite signs coexist in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer: the positive contribution for $t_{\mathrm{W}}<1$ nm and the negative contribution for $t_{\mathrm{W}}>1$ nm.
+The orbital-to-charge conversion with the negative sign for $t_{\mathrm{W}}>1$ nm can be attributed to the bulk inverse OHE in the W layer.
+The nature of the orbital-to-charge conversion can differ between the bulk and interfaces, accounting for the two orbital-to-charge conversion mechanisms with the opposite signs observed in the W($t_{\mathrm{W}}$)/Ni(5 nm) bilayer for $t_{\mathrm{W}}>1$ nm and $t_{\mathrm{W}}<1$ nm.
+Since the interface contribution originates from the anisotropic crystal potential, which differs from the bulk crystal potential go2024local , we assume that the interface contribution is particularly significant at the SiO2/W interface rather than at the metallic W/Ni interface supplementary .
 This interface effect should be especially strong at small $t_{\mathrm{W}}$, suggesting that the charge-current generation with the positive sign at $t_{\mathrm{W}}<1$ nm can be attributed to the orbital-to-charge conversion at the SiO2/W interface.
@@
 This indicates that the bulk inverse OHE dominates the observed orbital-to-charge conversion with the negative sign at larger $t_{\mathrm{W}}$.
-These results demonstrate that while the sign of the bulk direct OHE is positive, as evidenced by the torque measurements, the pumping measurements show that the sign of the bulk inverse OHE is negative.
-The observation of opposite signs for the bulk direct and inverse OHEs demonstrates that the local reciprocity is violated in the charge-orbital interconversion in the $\alpha$-W.
-Table 1: Decomposition of the bulk and surface contributions to the direct and inverse OHEs in W go2024local .
-The bulk and surface contributions to the direct(inverse) orbital Hall conductivity are represented as $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{OH}}^{\mathrm{dir(inv),surface}}$, respectively, along with the spin Hall conductivity components represented as $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),bulk}}$ and $\sigma_{\mathrm{SH}}^{\mathrm{dir(inv),surface}}$.
-Orbital/Spin Conversion Bulk Surface
-OHE Direct $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}<0$
-Inverse $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{OH}}^{\mathrm{inv,surface}}>0$
-SHE Direct $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{dir,bulk}}\simeq 0$
-Inverse $\sigma_{\mathrm{SH}}^{\mathrm{inv,bulk}}<0$ $\sigma_{\mathrm{SH}}^{\mathrm{inv,surface}}\simeq 0$
-Our observations are consistent with the theoretical prediction of a violation of local reciprocity between the direct and inverse OHEs in $\alpha$-W go2024local .
-In the theory, the non-reciprocity in the charge-orbital interconversion arises from the non-conservation of the OAM due to its strong interaction with the lattice via crystal fields.
-This interaction is accounted for by defining the orbital current as the sum of the conventional orbital current and the torque dipole, the latter being associated with the interaction between the OAM and the crystal lattice.
-It has been shown that the torque dipole is particularly significant at interfaces, where the anisotropic crystal potential efficiently mediates angular momentum exchange between the electrons and the lattice, strongly influencing the local profile of the charge-orbital interconversion in the W layer.
-As a result, the local charge currents induced by the orbital current deviate from the expectation of strict reciprocity.
-Notably, even the signs of the local responses differ between the direct and inverse OHEs, as summarized in Table I.
-The positive and negative signs of the bulk direct and inverse orbital Hall conductivities, $\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$ and $\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$, evidenced by the torque and pumping results, are consistent with this prediction.
-Importantly, these opposite signs do not violate Onsager’s reciprocity relation, as the direct and inverse responses are reciprocal when summed over the entire film thickness go2024local .
-These behaviors contrasts sharply with the SHE in $\alpha$-W, where the direct and inverse SHEs are largely reciprocal even at a local level go2024local .
-This indicates that the observed violation of the local reciprocity between the direct and inverse OHEs highlights the distinct nature of the charge-orbital interconversion compared to the charge-spin interconversion.
-Table I also suggests that the surface direct OHE with a negative sign may contribute to the negative torque for $t_{\mathrm{W}}<4$ nm.
+The above pumping results indicate that the bulk inverse OHE has the negative sign ($\sigma_{\mathrm{OH}}^{\mathrm{inv,bulk}}<0$), which is opposite to the positive sign of the bulk direct OHE ($\sigma_{\mathrm{OH}}^{\mathrm{dir,bulk}}>0$), as evidenced by the torque measurements.
+These opposite signs clearly demonstrate a violation of local reciprocity, as expressed by Eq.
+(2), in the charge-orbital interconversion in $\alpha$-W.
+The signs of the bulk direct and inverse OHEs, obtained directly from our thickness-dependent torque and pumping measurements, are consistent with the theoretical prediction summarized in Table I in Appendix A.
+Table I also suggests that the surface direct OHE with a negative sign may contribute to the negative torque for $t_{\mathrm{W}}<4$ nm, although its sign cannot be determined solely from the experimental data supplementary .
+Importantly, even though interfacial contributions cannot be fully quantified, the bulk contributions are reliably identified from the thickness dependence of the torque and pumping signals.
+Therefore, the key conclusion of this work—the direct and inverse OHEs in the bulk exhibit the opposite signs—remains robust.
+The observed local nonreciprocity in the OHEs contrasts sharply with the SHEs, where the direct and inverse SHEs remain largely reciprocal even at a local level in $\alpha$-W go2024local .
+This contrast highlights the distinct nature of charge-orbital interconversion compared with charge-spin interconversion.
 In summary, we have experimentally demonstrated the violation of local reciprocity in the charge-orbital interconversion in the W/Ni bilayers.
 By measuring the orbital torque and orbital pumping, we found that the sign of the bulk inverse OHE is opposite to that of the bulk direct OHE.
-This result confirms that the charge-orbital interconversion is locally non-reciprocal, in sharp contrast to the global reciprocity governed by Onsager’s reciprocal relations.
-The observation of the local reciprocity violation in the direct and inverse OHEs highlights the unique characteristics of charge-orbital coupled transport, providing essential insights for a deeper understanding of orbitronic phenomena.
+This result confirms that the charge-orbital interconversion is locally nonreciprocal, $\sigma_{\mathrm{OH}}^{\mathrm{dir}}(z)\neq\sigma_{\mathrm{OH}}^{\mathrm{inv}}(z)$, in sharp contrast to the global reciprocity expressed in Eq.
+(1), which is governed by Onsager’s reciprocal relations.
+The local violation of reciprocity between the direct and inverse OHEs is also expected in other heavy and light metals supplementary , suggesting that local nonreciprocity could serve as a powerful experimental signature of the OHE across a wide range of materials.
+Because local nonreciprocity originates from the nonconservation of OAM, extending such studies to systems where the loss of OAM can be controlled through interactions with magnons or phonons could open new directions for manipulating charge-orbital coupled transport.
+In spintronics, while the direct and inverse SHEs have generally been regarded as reciprocal processes, recent studies have revealed that nonreciprocal charge-spin conversion can emerge in Rashba systems or through spin-vorticity coupling PhysRevLett.122.217701 ; PhysRevB.96.140409 .
+These findings highlight that investigating reciprocity in charge-spin and charge-orbital interconversions is essential for achieving a unified understanding of angular-momentum transport in solids.
 Acknowledgements.
-This work was supported by JSPS KAKENHI (Grant Number: 22H04964, 23K19037), Spintronics Research Network of Japan (Spin-RNJ), and MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) (Grant Number: JPJ011438).
+We thank J.
+Yoon and T.
+Gao for valuable discussions, and K.
+Yamanoi and Y.
+Nozaki for their assistance in sample preparation.
+This work was supported by JSPS KAKENHI (Grant Numbers: 22H04964, 23K19037, 25K21707), Spintronics Research Network of Japan (Spin-RNJ), and MEXT Initiative to Establish Next-generation Novel Integrated Circuits Centers (X-NICS) (Grant Number: JPJ011438).
 Y.M.